what are two consecutive odd integers, such that seven times the smaller minus two times the larger is 201?
Answer by rapaljer(4671)

Let x = smaller integer
x+2 = second odd integer (larger)

7x- 2(x+2) =201
7x -2x-4 =201
5x =201+4
5x=205
x=41 smaller
x+2 = 43 larger

Check: 7*41 - 2*43 = 201??
287-86=201 Yes! It checks!!

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus